# Enhancing Fluorescence Lifetime Parameter Estimation Accuracy with Differential Transformer Based Deep Learning Model Incorporating Pixelwise Instrument Response Function

### v1

@@
Fluorescence lifetime imaging (FLI) is an important molecular imaging modality that can provide unique information for biomedical applications. FLI is based on acquiring and processing photon time of arrival histograms. The shape and temporal offset of these histograms depends on many factors, such as the instrument response function (IRF), optical properties, and the topographic profile of the sample. Several inverse solver analytical methods have been developed to compute the underlying fluorescence lifetime parameters, but most of them are computationally expensive and time-consuming. Thus, deep learning (DL) algorithms have progressively replaced computation methods in fluorescence lifetime parameter estimation. Often, DL models are trained with simple datasets either generated through simulation or a simple experiment where the fluorophore surface profile is mostly flat; therefore, DL models often do not perform well on samples with complex surface profiles such as ex-vivo organs or in-vivo whole intact animals. Herein, we introduce a new DL architecture using state-of-the-art Differential Transformer encoder-decoder architecture, MFliNet (Macroscopic FLI Network), that takes an additional input of IRF together with TPSF, addressing discrepancies in the photon time-of-arrival distribution. We demonstrate the model’s performance through carefully designed, complex tissue-mimicking phantoms and preclinical in-vivo cancer xenograft experiments.
@@ 1 Introduction
To address these challenges, understanding the underlying methodology for estimating lifetime parameters becomes important. In mFLI and MFLI, lifetime parameters can be estimated by deconvolving the temporal point spread function (TPSF) and instrument response function (IRF). TPSF is the temporal histogram of the acquired fluorescence photons exiting the surface of the sample after a pulse excitation. The IRF represents the temporal response of the imaging system to pulsed illumination [23]. Considering the complexity involved in estimating FLI parameters across diverse imaging conditions, fast and advanced data processing techniques are necessary to enhance both the precision and efficiency of these analyses. Recently, the field has seen a shift toward rapid, fit-free deep learning (DL) methodologies to alleviate the computational burden and reliance on user expertise, typically associated with methods such as nonlinear least squares fitting (NLSF) for FLI parameter estimation [24, 25]. This advancement makes real-time FLI a possibility [26, 27], driven by the development of novel DL methods that eliminate traditional time-consuming computational approaches. FLI-Net, a DL model developed for FLI parameter estimation [28], is used to analyze FLI data quickly, producing 2D quantitative images of the lifetimes and corresponding parameters directly without requiring manual parameter adjustments, outputting 2D quantitative images of the lifetime parameters directly. FLI-Net is versatile of FLI experiments, including visible and near-infrared (NIR) imaging, making it adaptable to a large range of biomedical applications. FLI-Net takes the TPSF as an input and outputs FLI parameters. The experimental IRF was used in data generation of the training data; hence, it was represented in the TPSFs. However, FLI-Net was not designed to analyze pixel-wise IRFs while it predicts the lifetime parameters.
@@
Following the developments in DL models, we leverage herein the ability of transformers to handle sequential data. Generally, transformers have a natural ability to capture long-range connections within data [29, 30, 31, 32]. In the context of FLI, they can effectively identify and learn the relationships between the TPSF and IRF for accurate lifetime estimation. Furthermore, the self-attention mechanism in transformers allows them to focus on the most relevant parts of the input data for making predictions [32]. In this study, we developed MFliNet to incorporate the Differential Transformer (DIFF Transformer) [33], enhancing its ability to process sequential data for accurate fluorescence lifetime parameter estimation. Our work introduces a novel decoder layer design that integrates the DIFF Transformer for the first time in literature, improving the model’s adaptability to account for shifts caused by variations in the DOF. The DIFF Transformer employs a differential attention mechanism, which calculates attention scores as the difference between two separate softmax attention maps, effectively canceling noise. This cancellation mechanism encourages sparse attention patterns, intensifying the focus on relevant contextual data while reducing distractions caused by irrelevant input. Analogous to noise-canceling systems, this approach mitigates the problem of over-allocating attention to non-critical information, a common issue with traditional transformers. DIFF Transformers outperform conventional transformers across various domains, especially in long-context modeling, key information retrieval, and robustness against variability in input structure. In the context of Time-Resolved FLI, these characteristics enable the detection of critical patterns even in low signal-to-noise scenarios, achieving higher accuracy and robustness compared to standard transformer architectures. These advancements mark a significant progression in FLI methodologies, offering a more effective tool for handling complex biological and imaging variability.
@@ 2.1 Imaging setup
All experimental data used in this work were captured on our MFLI system, where the detailed information can be found in [34]. Briefly, the system uses a large-format Intensified Charge-Coupled Device (ICCD) camera (Picostar HR, LaVision GmbH, Germany), for wide-field detection over a $8\times 6$ cm2 in combination with a Digital micro-mirrors device (DMD), DLi 4110, Texas Instruments, TX, USA, for a wide-field illumination. As an excitation source, we used a tunable Ti-Sapphire laser, Mai Tai HP, Spectra-Physics, CA, USA, which delivers $100$ femtosecond pulses at $80$ MHz. A gate width of 300 picoseconds (ps) and gate delay of $40$ ps were used for capturing time-resolved fluorescence decays (for in vivo and in vitro experiments) with a total of 176 time points, which is referred to as a number of gates (G=176). An emission filter at $740\times 10$ nm (FF01-740/13-25, Semrock, IL, Rochester, NY, USA) is used to capture the TPSFs, with the laser set at a $700$ nm wavelength, and Alexa Fluor 700 dye was used to obtain the time-resolved fluorescence signals.
@@ 2.2 Generation of training data and classical Fluorescence lifetime processing
The in silico data used for training and validating the proposed model was generated using Eq. 1. Initially, time-resolved fluorescence lifetime images with dimensions of $28\times 28$ pixels were generated by using the MNIST dataset. Fluorescence decays were generated for a range of lifetime values commonly used in NIR applications: $0.2$ ns to $0.8$ ns for $\tau_{1}$ (short-lifetime component) and $0.8$ ns to $1.5$ ns for $\tau_{2}$ (long-lifetime component). The range of the $A_{R}$ (fraction amplitude) was set from 0% to 100%, respectively (both bound corresponding to mono-exponential cases). To ensure that our simulated data accurately represents experimental applications, pixel-wise IRFs were used. To capture experimental IRFs, a white diffused paper was placed on the imaging table and illuminated using the DMD with an excitation wavelength of 700 nm. The reflected light was captured using a neutral density (ND) filter. Subsequently, each TPSF was generated by convolving randomly selected IRF from the dataset with simulated fluorescence decay profiles. To approximate the noise characteristics of real-world measurements, system-derived noise, including read-out noise, dark noise, etc., as explained in [34], was incorporated into the simulated TPSFs. This approach ensures that the simulated data closely matches the noise dynamics observed in the actual system.
To evaluate and compare the model’s performance in the absence of experimental ground truth, we used the NLSF method, which is commonly used to estimate the FLI parameters described in Eq. 1. Traditionally, FLI parameters are estimated from experimental data through iterative fitting optimization methods such as the NLSF, which incorporates the Levenberg-Marquardt algorithm [35], or center of mass (CMM) analysis [36]. For our NLSF analysis, we utilized a software named AlliGator [37], allowing adjustments and constraints on fitting parameters, including short and long lifetimes, fraction amplitudes, and offsets, depending on experimental conditions. We selected between single and double exponential decay models according to the complexity of our datasets. AlliGator also provides an option for offset correction when there is a mismatch between the TPSF and IRF. We evaluated the importance of offset correction in our NLSF analysis by comparing data with and without this feature in our phantom experiments. Additionally, we benchmarked our results against those obtained using FLI-Net to provide a comprehensive evaluation of our approach in the context of established methodologies.
@@ 2.3 Deep learning network architecture
MFliNet is a novel architecture designed for FLI parameter estimation, particularly effective under varying IRFs. The model leverages a Differential Transformer framework, incorporating a unique differential attention mechanism to enhance feature extraction while mitigating the effects of noise. The theoretical background on the differential attention mechanism is detailed in [33].
@@ 2.4 Phantom preparation
For experimental validation, we designed a step ladder phantom to introduce variations in sample-detector distance, as depicted in Figure 2(a). A 3D printable case was designed to accommodate five discrete containers arranged at various heights: ground level, $5$ mm, $10$ mm, $15$ mm, and $20$ mm. Each container was crafted with dimensions of $40\times 40\times 10$ mm to accommodate tissue-mimicking phantoms. The phantoms were made with agar constituting 1% of the total volume ($80$ cm3). To prepare the phantoms, agar was first dissolved in distilled water, heated until fully integrated, and then allowed to cool slightly before further processing. The optical properties of the phantoms (absorption coefficient ($\mu_{a}$) of $0.005$ mm-1 and a reduced scattering coefficient ($\mu_{s}^{\prime}$) of $1$ mm-1) were controlled through the addition of India Ink and intralipid solutions, to provide absorption and scattering contrasts respectively [38]. In each ladder step, a specific area was designated for the placement of a cuboidal fluorescence embedding, with dimensions of $5\times 5\times 40$ mm. This embedding consisted of Alexa Fluor 700 dye dissolved in phosphate-buffered saline to achieve a concentration of $20$ $\mu$M. The embeddings were placed at a depth of $1$ mm from the surface of each phantom.
Figure 2: Illustration of designed 3D step ladder phantom and the IRF shifts as a result of variations in height: (a) display of the 3D phantom (left) and plots of randomly selected pixelwise IRFs from each height (right); (b) a side view of a mouse, highlighting height differences between anatomical regions (left) and IRF plots of randomly selected pixels on liver, urinary bladder (UB), tumors (right); (c) A distal ventral view of the mouse highlighting the tumors and the liver (left) and the IRF plots of the randomly selected pixels on the left tumor (right).
@@
2.5 In vivo experiment
For in vivo MFLI imaging experiments, we imaged HER2+ breast tumor xenografts HCC1954 in athymic nude mice. The cell line was sourced from ATCC (Manassas, VA, USA) and maintained in RPMI 1640 media enriched with 10% fetal bovine serum (ATCC) and $50$ units/mL/ $50$ µg/mL penicillin/streptomycin from ThermoFisher Scientific (Waltham, MA, USA). We initiated tumor xenografts by subcutaneously injecting $5\times 10^{6}$ HCC1954 cells suspended in PBS and mixed in a $1:1$ ratio with Cultrex BME (R&D Systems Inc, Minneapolis, MN, USA) into the inguinal mammary fat pads of female athymic nude mice aged 4 weeks (CrTac: NCR-Foxn1nu, Taconic Biosciences, Rensselaer, NY, USA). Tumors were monitored daily for 4 weeks. The mouse was administered with a retro-orbital injection of AF700 conjugated with MDT-TZM (MDT-TZM-AF700) at $20$ µg and AF750 conjugated with MDT-TZM (MDT-TZM-AF750) at $40$ µg in a $2:1$ acceptor to donor ratio through staggered injection [9]. Donor injection was performed 2 hours ahead of acceptor injection through the retro-orbital route. MFLI Imaging was conducted 24 hours post-injection using the MFLI imaging setup. Throughout the imaging process, the mouse was anesthetized with isoflurane, and the body temperature was maintained with a Rodent Warmer X2 (Stoelting, IL, USA). All animal procedures were conducted with the approval of the Institutional Animal Care and Use Committee (IACUC) at both Rensselaer Polytechnic Institute and Albany Medical College. The animal facilities of both institutions have been accredited by the American Association for Accreditation for Laboratory Animals Care International.
@@ 3 Results
Figure 3: Phantom experiment results. a) Image overlay of the lifetime estimation results, b) Violin plots of NLSF analysis, FLI-Net, and MFliNet
We conducted the ladder phantom experiment designed to validate the MFliNet model under controlled conditions that mimic biological tissues. Figure 3 shows the phantom experiment results where the analysis was done using three methods: NLSF, FLI-Net, and MFliNet. To compare the precision and stability of each method under varying conditions reflective of real-world applications, results were evaluated across five different heights: ground, $5$ mm, $10$ mm, $15$ mm, and $20$ mm. A $160$ ps shift in the IRF was observed from ground level to a height of $20$ mm, with a shift of approximately $40$ ps for each $5$ mm increment in height. For simplicity in comparison, the amplitude-weighted average lifetime was calculated using Eq. 17, for all outputs.
@@
NLSF analysis using pixel-wise IRF showed consistency in lifetime estimation across all tested heights. The mean fluorescence lifetime values obtained by NLSF were clustered around $1.01\pm 0.02$ ns. NLSF without offset correction results deteriorated with each increase in height. At the ground level, NLSF without offset correction began with a mean value of $1.04\pm 0.01$ ns. However, as the height increased, a steady decline in lifetime estimation was observed, reaching a mean value of $0.93\pm 0.01$ ns at $20$ mm. FLI-Net, in contrast, demonstrated a wider variation in estimated fluorescence lifetime values. At the ground level, it reported a mean value of $0.96\pm 0.01$ ns, which was lower than the NLSF values. As the distance increased, FLI-Net’s estimations deviated further, peaking at $1.14\pm 0.01$ ns at $10$ mm and estimating $20$ mm with a mean value of $1.10\pm 0.02$ ns. In contrast, MFliNet, showed closer results with NLSF, where the mean values were within the same range as NLSF. Moreover, in terms of processing speed, NLSF took approximately 6 hours to analyze 598 pixels (covering only the tumor area), whereas MFliNet analyzed the entire dataset of 90,480 pixels in just 63 seconds.
@@
Following the phantom studies, in-vivo experiments were conducted using the HER2+ breast tumor xenograft model in mice to evaluate the model’s performance in a more complex, biologically variable environment. The experimental results, illustrated in Figure 4, demonstrate the comparative analysis of the HCC1954 cell line using NLSF and MFliNet. For the smaller tumor (HCC1954 (A)), the NLSF method reported a short fluorescence lifetime of $0.56$ ns with a standard deviation of $0.06$ ns and a long lifetime of $1.17\pm 0.03$ ns. The MFliNet showed a comparable short lifetime of $0.52\pm 0.05$ ns and a long lifetime of $1.19\pm 0.02$ ns. In the case of the larger tumor (HCC1954 (B)),the NLSF method reported a short fluorescence lifetime of $0.55$ ns with a standard deviation of $0.07$ ns and MFliNet showed a comparable short lifetime of $0.58\pm 0.04$ ns. For the long lifetime, both methods again yielded closely aligned values: $1.16\pm 0.04$ ns for NLSF and $1.18\pm 0.03$ ns for MFliNet.

### v2

@@
Fluorescence Lifetime Imaging (FLI) is a critical molecular imaging modality that provides unique information about the tissue microenvironment, which is invaluable for biomedical applications. FLI operates by acquiring and analyzing photon time-of-arrival histograms to extract quantitative parameters associated with temporal fluorescence decay. These histograms are influenced by the intrinsic properties of the fluorophore, instrument parameters, time-of-flight distributions associated with pixel-wise variations in the topographic and optical characteristics of the sample. Recent advancements in Deep Learning (DL) have enabled improved fluorescence lifetime parameter estimation. However, existing models are primarily designed for planar surface samples, limiting their applicability in translational scenarios involving complex surface profiles, such as in-vivo whole-animal or imaged guided surgical applications. To address this limitation, we present MFliNet (Macroscopic FLI Network), a novel DL architecture that integrates the Instrument Response Function (IRF) as an additional input alongside experimental photon time-of-arrival histograms. Leveraging the capabilities of a Differential Transformer encoder-decoder architecture, MFliNet effectively focuses on critical input features, such as variations in photon time-of-arrival distributions. We evaluate MFliNet using rigorously designed tissue-mimicking phantoms and preclinical in-vivo cancer xenograft models. Our results demonstrate the model’s robustness and suitability for complex macroscopic FLI applications, offering new opportunities for advanced biomedical imaging in diverse and challenging settings.
@@ 1 Introduction
To address these challenges, understanding the underlying methodology for estimating lifetime parameters becomes important. In mFLI and MFLI, lifetime parameters can be estimated by deconvolving the temporal point spread function (TPSF) and instrument response function (IRF). TPSF is the temporal histogram of the acquired fluorescence photons exiting the surface of the sample after a pulse excitation. The IRF represents the temporal response of the imaging system to pulsed illumination [23]. Considering the complexity involved in estimating FLI parameters across diverse imaging conditions, fast and advanced data processing techniques are necessary to enhance both the precision and efficiency of these analyses. Recently, the field has seen a shift toward rapid, fit-free deep learning (DL) methodologies to alleviate the computational burden and reliance on user expertise, typically associated with methods such as nonlinear least squares fitting (NLSF) for FLI parameter estimation [24, 25]. This advancement makes real-time FLI a possibility [26, 27], driven by the development of novel DL methods that eliminate traditional time-consuming computational approaches. FLI-Net, a DL model developed for FLI parameter estimation [28], is used to analyze FLI data quickly, producing 2D quantitative images of the lifetimes and corresponding parameters directly without requiring manual parameter adjustments, outputting 2D quantitative images of the lifetime parameters directly. FLI-Net is versatile in terms of the imaging domain, including visible and near-infrared (NIR) imaging, making it adaptable to a large range of biomedical applications. FLI-Net takes the TPSF as an input and outputs FLI parameters. The experimental IRF was used in data generation of the training data; hence, it was represented in the TPSFs. However, FLI-Net was not designed to analyze pixel-wise IRFs while it predicts the lifetime parameters.
@@
Following the developments in DL models, we leverage herein the ability of transformers to handle sequential data. Generally, transformers have a natural ability to capture long-range connections within data [29, 30, 31, 32]. In the context of FLI, they can effectively identify and learn the relationships between the TPSF and IRF for accurate lifetime estimation. Furthermore, the self-attention mechanism in transformers allows them to focus on the most relevant parts of the input data for making predictions [32]. Recently, Differential Transformer (DIFF Transformer) as been proposed as a new approach to amplify attention to the relevant task while canceling noise [33]. MFliNet architecture is based on a novel decoder layer design that integrates the DIFF Transformer for the first time in literature, improving the model’s adaptability to account for shifts caused by variations in the DOF. The DIFF Transformer employs a differential attention mechanism, which calculates attention scores as the difference between two separate softmax attention maps, effectively canceling noise. This cancellation mechanism encourages sparse attention patterns, intensifying the focus on relevant contextual data while reducing distractions caused by irrelevant input. Analogous to noise-canceling systems, this approach mitigates the problem of over-allocating attention to non-critical information, a common issue with traditional transformers. DIFF Transformers outperform conventional transformers across various domains, especially in long-context modeling, key information retrieval, and robustness against variability in input structure. In the context of Time-Resolved FLI, these characteristics enable the detection of critical patterns even in low signal-to-noise scenarios, achieving higher accuracy and robustness compared to standard transformer architectures. These advancements mark a significant progression in FLI methodologies, offering a more effective tool for handling complex biological and imaging variability.
@@ 2.1 Imaging setup
All experimental data used in this work were captured on our MFLI system, detailed information about which can be found in [34]. Briefly, the system uses a large-format Intensified Charge-Coupled Device (ICCD) camera (Picostar HR, LaVision GmbH, Germany), for wide-field detection over a $8\times 6$ cm2 in combination with a Digital micro-mirrors device (DMD) (DLi 4110, Texas Instruments, TX, USA), for wide-field illumination. As an excitation source, we used a tunable Ti-Sapphire laser (Mai Tai HP, Spectra-Physics, CA, USA), which delivers $100$ femtosecond pulses at $80$ MHz. A gate width of 300 picoseconds (ps) and gate delay of $40$ ps were used for capturing time-resolved fluorescence decays (for in-vivo and in-vitro experiments) with a total of 176 time points, which is referred to as a number of gates (G=176). An emission filter at $740\times 10$ nm (FF01-740/13-25, Semrock, IL, Rochester, NY, USA) is used to capture the TPSFs, with the laser set at a $700$ nm wavelength.
@@ 2.2 Generation of training data and classical Fluorescence lifetime processing
The in-silico data used for training and validating the proposed model was generated using Eq. 1. Initially, time-resolved fluorescence lifetime images with dimensions of $28\times 28$ pixels were generated by using the MNIST dataset. Fluorescence decays were generated for a range of lifetime values commonly used in NIR applications: $0.2$ ns to $0.8$ ns for $\tau_{1}$ (short-lifetime component) and $0.8$ ns to $1.5$ ns for $\tau_{2}$ (long-lifetime component). The range of the $A_{R}$ (fraction amplitude) was set from 0% to 100%, respectively (both bounds corresponding to mono-exponential cases). To ensure that our simulated data accurately represents experimental applications, pixel-wise IRFs were used. To capture experimental IRFs, a white diffused paper was placed on the imaging table and illuminated using the DMD with an excitation wavelength of 700 nm. The reflected light was captured using a neutral density (ND) filter. Subsequently, each TPSF was generated by convolving randomly selected IRF from the dataset with simulated fluorescence decay profiles. To approximate the noise characteristics of real-world measurements, system-derived noise, including read-out noise, dark noise, etc., as explained in [34], was incorporated into the simulated TPSFs. This approach ensures that the simulated data closely matches the noise dynamics observed in the actual system.
To evaluate and compare the model’s performance in the absence of experimental ground truth, we used the NLSF method, which is commonly used to estimate the FLI parameters described in Eq. 1. Traditionally, FLI parameters are estimated from experimental data through iterative fitting optimization methods such as the NLSF, which incorporates the Levenberg-Marquardt algorithm [35], or center of mass (CMM) analysis [36]. For our NLSF analysis, we utilized a software named AlliGator [37], allowing adjustments and constraints on fitting parameters, including short and long lifetimes, fraction amplitudes, and offsets, depending on experimental conditions. We selected between single and double exponential decay models according to the complexity of our datasets. AlliGator also provides an option for offset correction when there is a mismatch between the TPSF and IRF. We evaluated the importance of offset correction in our NLSF analysis by comparing data with and without this feature in our phantom experiments. Additionally, we benchmarked our results against those obtained using FLI-Net and to provide a comprehensive evaluation of our approach in the context of established methodologies. Furthermore, to assess the impact of the DIFF Transformer model, we compared our results with those from a transformer model of same architecture and parameters, trained on the same dataset.
@@ 2.3 Deep learning network architecture
MFliNet is a novel architecture designed for FLI parameter estimation, particularly effective under varying IRFs. The model leverages a DIFF Transformer framework, incorporating a unique differential attention mechanism to enhance feature extraction while mitigating the effects of noise. The theoretical background on the differential attention mechanism is detailed in [33].
@@ 2.4 Phantom preparation
For experimental validation, we designed a step ladder phantom to introduce variations in sample-detector distance, as depicted in Figure 2(a). A 3D printable case was designed to accommodate five discrete containers arranged at various heights: ground level, $5$ mm, $10$ mm, $15$ mm, and $20$ mm. Each container was crafted with dimensions of $40\times 40\times 10$ mm3 to accommodate tissue-mimicking phantoms. The phantoms were made with agar constituting 1% of the total volume ($80$ cm3). To prepare the phantoms, agar was first fully dissolved in distilled water by heating, and then allowed to cool slightly before further processing. The optical properties of the phantoms (absorption coefficient ($\mu_{a}$) of $0.005$ mm-1 and reduced scattering coefficient ($\mu_{s}^{\prime}$) of $1$ mm-1) were controlled through the addition of India Ink and intralipid solutions, to provide absorption and scattering contrasts respectively [38]. In each ladder step, a specific area was designated for the placement of a cuboidal fluorescence embedding, with dimensions of $5\times 5\times 40$ mm3. This embedding consisted of Alexa Fluor 700 dye dissolved in phosphate-buffered saline to achieve a concentration of $20$ $\mu$M. The embeddings were placed at a depth of $1$ mm from the surface of each phantom.
Figure 2: Illustration of designed 3D step ladder phantom and the IRF shifts as a result of variations in height: (a) display of the 3D phantom (up) and plots of average of IRFs from each height (down); (b) a side view of a mouse, highlighting height differences between anatomical regions (up) and IRF plots of randomly selected pixels on liver, urinary bladder (UB), tumors (down); (c) A distal ventral view of the mouse highlighting the tumors and the liver (up) and the IRF plots of the randomly selected pixels on the left tumor (down).
@@
2.5 In-vivo experiment
For in-vivo MFLI imaging experiments, we imaged HER2+ breast tumor xenografts HCC1954 in athymic nude mice. The cell line was sourced from ATCC (Manassas, VA, USA) and maintained in RPMI 1640 media enriched with 10% fetal bovine serum (ATCC) and $50$ units/mL/ $50$ µg/mL penicillin/streptomycin from ThermoFisher Scientific (Waltham, MA, USA). We initiated tumor xenografts by subcutaneously injecting $5\times 10^{6}$ HCC1954 cells suspended in PBS and mixed in a $1:1$ ratio with Cultrex BME (R&D Systems Inc, Minneapolis, MN, USA) into the inguinal mammary fat pads of female athymic nude mice aged 4 weeks (CrTac: NCR-Foxn1nu, Taconic Biosciences, Rensselaer, NY, USA). Tumors were monitored daily for 4 weeks. The mouse was administered with a retro-orbital injection of AF700 conjugated with Meditope Trastuzumab (MDT-TZM) (MDT-TZM-AF700) at $20$ µg and AF750 conjugated with MDT-TZM (MDT-TZM-AF750) at $40$ µg in a $2:1$ acceptor to donor ratio through staggered injection [9]. Donor injection was performed 2 hours ahead of acceptor injection through the retro-orbital route. MFLI Imaging was conducted 24 hours post-injection. Throughout the imaging process, the mouse was anesthetized with isoflurane, and the body temperature was maintained with a Rodent Warmer X2 (Stoelting, IL, USA). All animal procedures were conducted with the approval of the Institutional Animal Care and Use Committee (IACUC) at both Rensselaer Polytechnic Institute and Albany Medical College. The animal facilities of both institutions have been accredited by the American Association for Accreditation for Laboratory Animals Care International.
@@ 3 Results
Figure 3: Phantom experiment results. a) Image overlay of the lifetime estimation results, b) Violin plots of NLSF analysis, FLI-Net, transformer model and MFliNet
We conducted the ladder phantom experiment designed to validate the MFliNet model under controlled conditions that mimic biological tissues. Figure 3 shows the phantom experiment results where the analysis was done using four methods: NLSF, FLI-Net, transformer model and MFliNet. To compare the precision and stability of each method under varying conditions reflective of real-world applications, results were evaluated across five different heights: ground, $5$ mm, $10$ mm, $15$ mm, and $20$ mm. A $160$ ps shift in the IRF was observed from ground level to a height of $20$ mm, with a shift of approximately $40$ ps for each $5$ mm increment in height. For simplicity in comparison, the amplitude-weighted average lifetime was calculated using Eq. 17, for all outputs.
@@
NLSF analysis using pixel-wise IRF showed consistency in lifetime estimation across all tested heights. The mean fluorescence lifetime values obtained by NLSF were clustered around $1.01\pm 0.02$ ns. NLSF without offset correction results deteriorated with each increase in height. At the ground level, NLSF without offset correction began with a mean value of $1.04\pm 0.01$ ns. However, as the height increased, a steady decline in lifetime estimation was observed, reaching a mean value of $0.93\pm 0.01$ ns at $20$ mm. FLI-Net, in contrast, demonstrated a wider variation in estimated fluorescence lifetime values. At the ground level, it reported a mean value of $0.96\pm 0.01$ ns, which was lower than the NLSF values. As the distance increased, FLI-Net’s estimations deviated further, peaking at $1.14\pm 0.01$ ns at $10$ mm and estimating $20$ mm with a mean value of $1.10\pm 0.02$ ns. The transformer model showed variable performance across different heights; at 10 mm, a decrease was observed with a mean value of $0.91\pm 0.04$ ns, while estimations at other heights were closer to the NLSF values. In contrast, MFliNet, showed closer results with NLSF, where the mean values were within the same range as NLSF. Moreover, in terms of processing speed, NLSF took approximately 6 hours to analyze 598 pixels (covering only the tumor area), whereas MFliNet analyzed the entire dataset of 90,480 pixels in just 63 seconds.
@@
Following the phantom studies, in-vivo experiments were conducted using the HER2+ breast tumor xenograft model in mice to evaluate the model’s performance in a more complex, biologically variable environment. The experimental results, illustrated in Figure 4, demonstrate the comparative analysis of NLSF and MFliNet in HER2+ HCC1954 tumor xenograft mice model. For the smaller tumor (HCC1954 (A)), the NLSF method reported a short fluorescence lifetime of $0.56\pm 0.06$ ns and a long lifetime of $1.17\pm 0.03$ ns. The MFliNet showed a comparable short lifetime of $0.52\pm 0.05$ ns and a long lifetime of $1.19\pm 0.02$ ns. In the case of the larger tumor (HCC1954 (B)), the NLSF method reported a short fluorescence lifetime of $0.55\pm 0.07$ ns and MFliNet showed a comparable short lifetime of $0.58\pm 0.04$ ns. For the long lifetime, both methods again yielded closely aligned values: $1.16\pm 0.04$ ns for NLSF and $1.18\pm 0.03$ ns for MFliNet.